HETAV PATEL

Grade 11,

The sum of the mole fraction of A,B,C in an aqueous solution containing.0.2 moles of each A,B and C is

The sum of the mole fraction of A,B,C in an aqueous solution containing.0.2 moles of each A,B and C is

Grade:12

3 Answers

Naureen Nida
5 years ago
given, nA = nB = nC  =  0.2
   mole fraction of A = nA÷nA+nB+nC =0.2÷0.2+0.2+0.2
                                = 0.2÷0.6=1/3
mole fraction of A =B=C= 1/3
Agrata Singh
5 years ago
Sum of mole fraction of all constituent of a mixture is always 1. No matter how many constituents are given and no matter their individual no of moles, if sum of all mole fractions is asked it is always one.
 
Aastha Singh
5 years ago
DEAR STUDENT,
MOLE FRACTION OF N NO. OF COMPOUNDS TOGETHER ALWAYS ADD UP TO 1
HERE moles of a,b and c = NA=NB=NC=0.2
MOLEFRACTIONS OF ALL THE THREE= 0.2/0.2+0.2+0.2= 1/3
SUM OF MOLE FRACTIONS OF ALL= 3(1/3)= 1